OpenStudy (anonymous):

The numerator of a fraction is 3 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 16/12. Find the original fraction.

OpenStudy (anonymous):

\[\text{Answer} \text{ is: } \frac{21}{7} \]Let d be the denominator. Then 3d is the numerator. Solve the following equation for d:\[\frac{3d-5}{d+5}\text{=}\frac{16}{12}\]

OpenStudy (anonymous):

\[d=7, 3d=21 \]